OpenStudy (anonymous):

The graph of f(x) = 2x + 4 shifts six units to the right when it is replaced with the graph of f(x) = 2x - k. What is the value of k? 2 4 6 10

OpenStudy (ranga):

Moving 6 units to the right is same as replacing x by (x-6) in the original function.

OpenStudy (ranga):

f(x) = 2x + 4 Shift 6 units to the right by replacing x by (x - 6) 2(x-6) + 4 = 2x - 12 + 4 = 2x - 8. So k = 8. None of the answer choices matches this answer but you can verify it is correct by plotting 2x + 4 and 2x - 8 and you will find the second graph is 6 units to the right of the first graph.